# Interference patterns for simple lens models in wave-optics regime

### v1

@@ Abstract
This work studies interference patterns created by simple lens models (point mass, Chang-Refsdal, and binary lens) in the wave optics regime, primarily in the context of lensing of gravitational waves (GWs) in the LIGO band at frequencies around $100$ Hz. We study how the interference patterns behave close to the caustic curves which mark the high magnification regions in conventional geometric optics. In addition, we also look at the formation of highly de-amplified regions in the amplification maps close to caustics and how they differ under wave and geometric optics. We see that for a source close to caustics, the oscillations in the amplification factor (their amplitude and location of crests and troughs) can differ significantly in wave optics compared to geometric optics. As we move away from caustics, the wave optics amplification factor starts to converge towards geometric optics one, especially the frequencies at which crests and through occur in the amplification factor, although the amplitude of these oscillations can still be considerably different. For Chang-Refsdal and binary lens with ${\sim}100\>{\rm M_{\odot}}-200\>{\rm M_{\odot}}$ can introduce significant de-amplification at frequencies ${\sim}100$ Hz when the source is close to caustics which may help us distinguish such lenses from the point mass lens.
@@ 1 Introduction
With the continuous detection of gravitational waves (GWs) by the LIGO-Virgo-KAGRA (LVK) network (e.g., Abbott et al., 2023, 2024) and the fact that gravitational lensing by stellar mass objects can introduce chromatic (i.e., frequency-dependent) effects in this frequency range, it has became crucial to understand gravitational lensing effects in the wave optics regime. Generally, wave effects become important if the lens mass is $M\lesssim 10^{5}(f/{\rm Hz})^{-1}$, where $f$ is the frequency of the incoming signal (e.g., Takahashi & Nakamura, 2003; Bulashenko & Ubach, 2022). More specifically, for GW signals in the LVK frequency range, the above mass primarily contains microlenses (i.e., stars and stellar mass compact object) and milli-lenses (i.e., star cluster and sub-halos).
@@
Under geometric optics approximation, a background GW source is multiply imaged, and for a given source position, the contribution to the amplification factor (also known as transmission factor; e.g., Born & Wolf, 1999) only comes from stationary points of the Fermat potential (e.g., Blandford & Narayan, 1986) which essentially denote positions of multiple images in the image plane. In generic lensing scenarios under geometric optics, the time delays between these multiple images are large (ranging from days to years) compared to the length of the GW signal (${\lesssim}1$ second) in the LVK band, and we expect to observe one image at a time, (de-)amplified and phase-shifted by a constant factor. Once we go into the wave optics regime, we do not observe these multiple-lensed images. Although, as we will see below, images formed in geometric optics (i.e., virtual images) are still relevant in understanding various features in the wave optics regime. In wave optics, the incoming wavefront is treated as a constant phase surface, and the lens (plane) can be thought of as a transparent phase screen introducing an extra phase (which depends on the frequency and lens parameters) for every partial wave of the incoming wavefront (e.g., Schneider et al., 1992; Born & Wolf, 1999; Bulashenko & Ubach, 2022), and the observer detects the resulting interference pattern, including contributions also from non-stationary points. Due to this extra contribution in wave optics, the resulting interference pattern is also sensitive to the lens model properties and can help break various lensing degeneracies (Meena, 2024a).
@@
Most of the previous works have focused on studying the wave effects assuming a fixed source (or observer) position and as a function of frequency, for example, for an isolated point mass (i.e., Schwarzschild) lens (e.g., Takahashi & Nakamura, 2003; Bulashenko & Ubach, 2022), from a point mass lens in the presence of external effects (e.g., Meena & Bagla, 2020; Mishra et al., 2021; Yeung et al., 2023), and from an ensemble of point mass lenses equivalent to microlens population in galaxy and galaxy cluster lenses (e.g., Diego et al., 2019; Mishra et al., 2021; Meena et al., 2022; Shan et al., 2023; Meena, 2024b). This can be understood from the fact that, realistically, an observed lensed GW signal would span a range of frequencies and have one source position. In addition, at present, to our knowledge, we do not have efficient methods to simulate a large number of amplification factors for different source positions [^1] Except for point mass lens for which we have an analytical formula for amplification factor.. However, this can limit us from visualising the effect of the variation of the source position on the amplification factor at a given frequency (or as a function of frequency), especially close to the caustics where extra virtual images (under the geometric optics approximation) appear or disappear.
@@
In our current work, we simulate interference patterns in the source plane for simple lens models (point mass, Chang-Refsdal, and binary lens), primarily focusing on GW signals in the LVK frequency range. For the point mass lens case, thanks to the circular symmetry, the source position can be given by its polar coordinate in the source plane, and an amplification map can be constructed in a two dimension plane of source position and dimensionless frequency (see Fig. 4 in Bulashenko & Ubach, 2022). However, the same cannot be done once we break the circular symmetry as we have three-dimensional space with two parameters describing the source position and one for (dimensionless) frequency. Hence, for simplicity, we choose three frequency values $f=50,100,500$ Hz to construct interference maps as they cover the range of frequencies where nearly all signals detected by the LVK network are expected to have non-zero powers. We also discuss how our results can be used for other frequencies and lens masses. With these simulated maps, we study how the interference map varies as the number of virtual images changes close to caustics. It is possible that for certain lens system configurations (or source positions if everything else is fixed), different parts of the incoming wavefront interfere destructively and lead to highly de-amplified regions. Under geometric optics, finding such configurations is easy, and it would be interesting to see how closely geometric optics follow the full wave optics solution. Hence, we also look at the formation of such regions in the source plane under wave optics and at what lens mass range and source positions they occur at frequencies ${\sim}100$ Hz so that we may detect them in LVK signals.
@@
Figure 1: Amplification factor, $F(f)$, for point mass lens. Top-row: Source plane maps of the absolute value of amplification factor ($|F|$) at three different frequencies. Bottom-left: $\tilde{F}(t)$ curves normalized by $2\pi$ for three different values of source positions. Bottom-middle: Absolute value of amplification factor ($|F|$). The black dashed curves represent $|F|$ calculated using the analytical formula in Equation (13), and coloured solid curves are calculated using the UG method. The dashed-dotted curves show the $|F|$ under geometric optics approximation. Bottom-right: Phase value of amplification factor ($\theta_{F}$). Like the middle panel, Black dashed curves are estimated using the analytic formula. Coloured solid curves are calculated using the UG method, and dashed-dotted curves are the results under the geometric optics approximation.
@@ 2 Lensing basics in wave optics regime
This section briefly revisits the relevant basics of gravitational lensing in wave optics regimes. For a more detailed description, we refer readers to the following works: Schneider et al. (1992); Nakamura & Deguchi (1999); Takahashi & Nakamura (2003).
@@
$F(f,\boldsymbol{y})=\frac{1+z_{d}}{\rm c}\frac{D_{d}D_{s}}{D_{ds}}\theta_{0}^{2}\frac{f}{\iota}\int d^{2}\boldsymbol{x}\exp[2\pi\iota ft_{d}(\boldsymbol{x},\boldsymbol{y})],$ (1)
where $z_{d}$ is the lens redshift. $\boldsymbol{x}\equiv\boldsymbol{\theta}/\theta_{0}$ and $\boldsymbol{y}\equiv\boldsymbol{\beta}/\theta_{0}$ are the dimensionless position vectors in the image and source plane, respectively, with $\theta_{0}$ being the normalization angular scale. $D_{d}$, $D_{s}$ and $D_{ds}$ are angular diameter distances from observe to the lens, observer to source, and lens to source, respectively. $t_{d}(\boldsymbol{x},\boldsymbol{y})$ represents the well arrival time delay surface given as,
@@
where $\psi(\boldsymbol{x})$ is the lensing potential and $\phi(\boldsymbol{y})$ is a constant independent of lens properties, which is usually set to a value such that the arrival time delay for global minima is zero. As we can see from the above equation, the amplification factor, $F(f)$, is a frequency-dependent complex-valued function. For GWs, the change in amplitude of the signal due to lensing is given by the corresponding absolute value $(|F|)$ and lensing-induced phase-shift is given by $\theta_{F}\equiv-\iota\ln[F/|F|]$. In geometric optics limit $(ft_{d}\gg 1)$, the integral in Equation (1) becomes highly oscillatory, and only its stationary points contribute to the amplification factor such that
$F(f,\boldsymbol{y})|_{\rm geo}=\sum_{j}\sqrt{|\mu_{j}|}\exp\left(2\pi\iota ft_{d}(\boldsymbol{x}_{j},\boldsymbol{y})-\iota\pi n_{j}\right),$ (3)
where $\mu_{j}$ is the geometric optics magnification of the $j$-th image and $n_{j}$ is the Morse index with values 0, 1/2, and 1 for minima, saddle-point, and maxima images, respectively (e.g., Schneider et al., 1992; Nakamura & Deguchi, 1999). The extra phase shift arises from complex Gaussian integral (close to stationary points) and can be used to identify lensed GW signals (e.g., Ezquiaga et al., 2021; Wang et al., 2021; Vijaykumar et al., 2023).
Often, at least for simple lens models discussed in current work, it is useful to write the time delay and frequency in dimensionless quantities, which can be defined as (Mishra et al., 2021),
@@
$\begin{split}F(\nu,\boldsymbol{y})&=\frac{\nu}{\iota}\int d^{2}\boldsymbol{x}\exp[2\pi\iota\nu\tau_{d}(\boldsymbol{x},\boldsymbol{y})],\\ F(\nu,\boldsymbol{y})|_{\rm geo}&=\sum_{j}\sqrt{|\mu_{j}|}\exp\left(2\pi\iota\nu\tau_{d}(\boldsymbol{x}_{j},\boldsymbol{y})-\iota\pi n_{j}\right).\end{split}$ (5)
As we will see below, such a dimensionless form allows us to translate results for one particular lens mass ($M_{z}$) or frequency ($f$) to another, at least for the lens models considered in our work.
@@ 3 Calculating $F(f)$
Except for very simple lens models, one cannot obtain an analytic solution for Equation (1) or (5). Hence, for a general lens, we need to rely on numerical integration or some other methods (e.g., Ulmer & Goodman, 1995; Diego et al., 2019) to determine the amplification factor. In our current work, we primarily use the method described in Ulmer & Goodman (1995, ‘UG method’ hereafter), which relies on the Fourier transformation and contour integration. First, assuming a constant source position, we take the Fourier transformation of $\iota F(\nu)/\nu$ in Equation (5) defined as,
[…]
@@
$\tilde{F}(\tau^{\prime})=\int d^{2}\boldsymbol{x}\>\delta[\tau(\boldsymbol{x})-\tau^{\prime}].$ (7)
@@
[…]
@@
Figure 2: Amplification factor, $|F|$, maps for Chang-Refsdal lens with positive values of external shear (i.e., $\gamma>0$). The three rows represent the source plane maps corresponding to $\gamma=0.5/\sqrt{2},0.8/\sqrt{2},0.9/\sqrt{2}$ (i.e., $\mu=1.33,2.78,5.26$), respectively. In each row, left, middle and right panels correspond to $f=50,100,500$ Hz. In each panel, black solid curves represent the caustic. The black and green stars in the top and middle rows represent the source position for which amplification factor curves are shown in Figure 4.
@@ 4 Point mass lens
For the point mass lens, the amplification factor (1) has an analytical solution, which is given as,
$\begin{split}F(f,y)=\exp\left\{\frac{\pi\omega}{4}+\frac{\iota\omega}{2}\left[\ln\left(\frac{\omega}{2}\right)-2\phi_{\rm m}(y)\right]\right\}\\ \Gamma\left(1-\frac{\iota\omega}{2}\right){}_{1}F_{1}\left(\frac{\iota\omega}{2},1;\frac{\iota\omega y^{2}}{2}\right),\end{split}$ (13)
@@
where $\omega\equiv 2\pi\nu$. $\phi_{\rm m}(y)=(x_{\rm m}-y)^{2}/2-\ln(x_{\rm m})$ with $x_{\rm m}=(y+\sqrt{y^{2}+4})/2$ represents the time delay corresponding to the global minima image in geometric optics such that it arrives at $t_{d}=0$. In geometric optics limit ($\nu\tau_{d}\gg 1$), the amplification factor can be written as,
$F(f,y)|_{\rm geo}=\sqrt{|\mu_{+}|}-\iota\sqrt{|\mu_{-}|}\exp[2\pi\iota\nu\tau_{d}(x,y)],$ (14)
@@
where $\mu_{+}$ and $\mu_{-}$ are the usual magnification factors of minima and saddle-point images, respectively.
@@
Gravitational lensing by a point mass lens can be described by two parameters: redshifted lens mass ($M_{z}$) and source position ($y$). Due to the axial symmetry in the case of point mass lens, we can expect the interference pattern to show rings of constructive and destructive interference with the distance between consecutive maxima and minima depending on the lens mass and incoming signal frequency (Bulashenko & Ubach, 2022). The amplification factors for a point mass lens with $M_{z}=200~{}{\rm M_{\odot}}$ are shown in Figure 1. The top row represents the source plane map of the amplification factor ($|F|$) for three different frequencies, $f=50,100,500$ Hz, and the bottom row represents the $F(t)$, $|F|$, and $\theta_{F}$ for three different source positions. These amplification maps can also represent other lens masses or frequencies for the same source position such that they lead to the same value of dimensionless frequency ($\nu=4{\rm G}M_{z}f/{\rm c^{3}}$).
@@
In the case of an isolated point mass lens, we always have the formation of two images in geometric optics, with the saddle-point image becoming less and less amplified (approaching zero), and the minima image amplification factor approaches one as the source moves away from the optical axis (i.e., origin). At the same time, the time delay between the two images also increases. Using the geometric optics approximation from Equation (14), we can see that the increase in time delay implies that oscillations in amplification factor would become more rapid at a given frequency, and we move towards geometric optics at relatively lower frequencies. The larger fractional difference in the amplification of two images decreases the amplitude of the oscillations in the amplification factor. Both of the above can be seen from the dashed-dotted curves in the bottom row of Figure 1. In addition, we can also see that the amplification factor under geometric optics approximation and wave optics mainly differs at frequencies lower than the first peak in the interference pattern, implying that above the frequency of the first peak, the contribution to the amplification factor is well localized around the stationary points and as we lower the frequency, the contribution from regions far away from stationary points increases. For the perfect alignment (i.e., $y=0$) case, we have the formation of a ring with zero time delay, implying that we reach geometric optics at $f=\infty$ and wave effects are important for every observable frequency.
[…]
@@
Figure 6: Example amplification factor, $F(f)$, for binary lens case. Source positions are shown by solid stars in the third and fourth rows of Figure 5. The left, middle, and right panels show the $F(t)$$|F|$, and $\theta_{F}$ curves, respectively. The solid curves are obtained with the UG method and dashed curves represent the amplification factors obtained under geometric optics. In the bottom panel, for comparison, the dashed-dotted curves represent the amplification factor curves for a point mass lens with $M_{Tz}=200~{}{\rm M_{\odot}}$ and $y=|\boldsymbol{y}|$.

### v2

@@ Abstract
This work studies interference patterns created by simple lens models (point mass, Chang-Refsdal, and binary lens) in the wave optics regime, primarily in the context of lensing of gravitational waves (GWs) in the LIGO band at frequencies around $100$ Hz. We study how the interference patterns behave close to the caustic curves, which mark the high magnification regions in conventional geometric optics. In addition, we also look at the formation of highly de-amplified regions in the amplification maps close to caustics and how they differ under wave and geometric optics. We see that, except for close to caustic regions, the geometric optics track the oscillations of the amplification factor in frequency very well, although the amplitude of these oscillations can differ considerably. Chang-Refsdal and binary lenses with masses ${\sim}100\>{\rm M_{\odot}}-200\>{\rm M_{\odot}}$ can introduce significant de-amplification at frequencies ${\sim}100$ Hz when the source is close to caustics, which may help us distinguish such lenses from the point mass lens.
keywords:
gravitational lensing – gravitational waves.
[…]
@@ 1 Introduction
With the continuous detection of gravitational waves (GWs) by the LIGO-Virgo-KAGRA (LVK) network (e.g., Abbott etalet~al., 2023, 2024) and the fact that gravitational lensing by stellar mass objects can introduce chromatic (i.e., frequency-dependent) effects in the corresponding frequency range, it has become crucial to understand gravitational lensing effects in the wave optics regime. Generally, wave effects become important if the lens mass is $M\lesssim 10^{5}(f/{\rm Hz})^{-1}$, where $f$ is the frequency of the incoming signal (e.g., Takahashi & Nakamura, 2003; Bulashenko & Ubach, 2022). More specifically, for GW signals in the LVK frequency range, the above mass primarily contains microlenses (e.g., stars and stellar mass compact objects) and milli-lenses (e.g., star clusters and sub-halos).
@@
Under geometric optics approximation, a background GW source is multiply imaged, and for a given source position, the contribution to the amplification factor (also known as transmission factor; e.g., Born & Wolf, 1999) only comes from stationary points of the Fermat potential (e.g., Blandford & Narayan, 1986), which essentially denote positions of multiple images in the image plane. In generic lensing scenarios under geometric optics, the time delays between these multiple images are large (ranging from days to years) compared to the length of the GW signal (${\lesssim}1$ second) in the LVK frequency band, and we expect to observe one image at a time, (de-)amplified (and phase-shifted) by a constant factor. Once we go into the wave optics regime, we do not observe these multiple lensed images. Although, as we will see below, images formed in geometric optics (i.e., virtual images) are still relevant in understanding various features in the wave optics regime. In wave optics, the incoming wavefront is treated as a constant phase surface, and the lens (plane) can be thought of as a transparent phase screen introducing an extra phase (which depends on the frequency and lens parameters) for every partial wave of the incoming wavefront (e.g., Schneider etalet~al., 1992; Born & Wolf, 1999; Bulashenko & Ubach, 2022), and the observer detects the resulting interference pattern, including contributions also from non-stationary points. Due to this extra contribution in wave optics, the resulting interference pattern is also sensitive to the lens model properties and can help break various lensing degeneracies (Tambalo etalet~al., 2023; Meena, 2024b; Poon etalet~al., 2025).
@@
Most of the previous works have focused on studying the wave effects assuming a fixed source (or observer) position and as a function of frequency for an isolated point mass (i.e., Schwarzschild) lens (e.g., Takahashi & Nakamura, 2003; Bulashenko & Ubach, 2022), from a point mass lens in the presence of external effects (e.g., Meena & Bagla, 2020; Mishra etalet~al., 2021; Yeung etalet~al., 2023), and from an ensemble of point mass lenses equivalent to microlens population in galaxy and galaxy cluster lenses (e.g., Diego etalet~al., 2019; Mishra etalet~al., 2021; Meena etalet~al., 2022; Shan etalet~al., 2023; Meena, 2024a). This can be understood from the fact that, realistically, an observed lensed GW signal would span a range of frequencies and have one source position. In addition, at present, to our knowledge, we do not have efficient methods to simulate a large number of amplification factors for different source positions [^1] Except for a point mass lens for which we have an analytical formula for the amplification factor.. This can limit us from visualising the effect of the variation of the source position on the amplification factor at a given frequency, especially close to the caustics, where extra virtual images (under the geometric optics approximation) appear or disappear.
@@
In our current work, we simulate interference patterns/maps in the source plane for simple lens models (point mass, Chang-Refsdal, and binary lens), primarily focusing on GW signals in the LVK frequency range. Strictly speaking, we simulate maps of the absolute value of the amplification factor. However, an interference pattern is essentially a map of $\rm|amplification|^{2}$ and we use the above terminologies interchangeably. For the point mass lens case, thanks to the circular symmetry, the source position can be given by its polar coordinate, and an amplification map can be constructed on the source position and dimensionless frequency plane, which can be re-scaled for a given lens mass (see Fig. 4 in Bulashenko & Ubach, 2022). However, the same cannot be done once we break the circular symmetry, as we have three-dimensional space with two parameters describing the source position and one for (dimensionless) frequency. Hence, for simplicity, we choose three frequency values $f=50,100,500$ Hz to show interference maps as they cover the range of frequencies where nearly all signals detected by the LVK network are expected to have non-zero powers. We also discuss how our results can be used for other frequencies and lens masses. With these simulated maps, we study how the interference map varies as the number of virtual images changes close to caustics. It is possible that for certain lens system configurations (or source positions assuming everything else is fixed), different parts of the incoming wavefront interfere destructively and lead to highly de-amplified regions. Under geometric optics, finding such configurations is easy, and it would be interesting to see how closely geometric optics follows the full wave optics solution. Hence, we also look at the formation of such regions in the source plane under wave optics and at what lens mass range and source positions they occur at frequencies ${\sim}100$ Hz so that we may detect them in LVK signals.
@@
Figure 1: Amplification factor, $F(f)$, for point mass lens. Top-row: Source plane maps of the absolute value of amplification factor ($|F|$) at three different frequencies. Bottom-left: $\tilde{F}(t)$ curves normalized by $2\pi$ for three different values of source positions marked by the black solid stars in the top row. Bottom-middle: Absolute value of amplification factor ($|F|$). The black dashed curves represent $|F|$ calculated using the analytical formula in Equation (13), and coloured solid curves are calculated using the UG method. The dashed-dotted curves show the $|F|$ under geometric optics approximation. Bottom-right: Phase value of amplification factor ($\theta_{F}$). Like the middle panel, the black dashed curves are estimated using the analytic formula. Coloured solid curves are calculated using the UG method, and dashed-dotted curves are the results under the geometric optics approximation.
@@ 2 Lensing basics in wave optics regime
This section briefly revisits the relevant basics of gravitational lensing in wave optics regimes. For a more detailed description, we refer readers to the following works: Schneider etalet~al. (1992); Nakamura & Deguchi (1999); Takahashi & Nakamura (2003).
@@
$F(f,\boldsymbol{y})=\frac{1+z_{d}}{\rm c}\frac{D_{d}D_{s}}{D_{ds}}\theta_{0}^{2}\frac{f}{i}\int d^{2}\boldsymbol{x}\exp[2\pi ift_{d}(\boldsymbol{x},\boldsymbol{y})],$ (1)
where $z_{d}$ is the lens redshift. $\boldsymbol{x}\equiv\boldsymbol{\theta}/\theta_{0}$ and $\boldsymbol{y}\equiv\boldsymbol{\beta}/\theta_{0}$ are the dimensionless position vectors in the image and source plane, respectively, with $\theta_{0}$ being the normalization angular scale. $D_{d}$, $D_{s}$, and $D_{ds}$ are angular diameter distances from observe to the lens, observer to source, and lens to source, respectively. $t_{d}(\boldsymbol{x},\boldsymbol{y})$ represents the well-known arrival time delay surface given as,
@@
where $\psi(\boldsymbol{x})$ is the lensing potential and $\phi_{m}(\boldsymbol{y})$ is a constant independent of lens properties, which is usually set to a value such that the arrival time delay for global minima is zero. As we can see from the above equation, the amplification factor, $F(f)$, is a frequency-dependent complex-valued function. For GWs, the change in amplitude of the signal due to lensing is given by the corresponding absolute value $(|F|)$ and lensing-induced phase-shift is given by $\theta_{F}\equiv-i\ln[F/|F|]$. In geometric optics limit $(ft_{d}\gg 1)$, the integral in Equation (1) becomes highly oscillatory, and only its stationary points contribute to the amplification factor such that
$F(f,\boldsymbol{y})|_{\rm geo}=\sum_{j}\sqrt{|\mu_{j}|}\exp\left(2\pi ift_{d}(\boldsymbol{x}_{j},\boldsymbol{y})-i\pi n_{j}\right),$ (3)
where $\mu_{j}$ is the geometric optics magnification of the $j$-th image and $n_{j}$ is the Morse index with values 0, 1/2, and 1 for minima, saddle-point, and maxima images, respectively (e.g., Schneider etalet~al., 1992; Nakamura & Deguchi, 1999). The extra phase shift arises from complex Gaussian integral (close to stationary points) and can be used to identify lensed GW signals (e.g., Ezquiaga etalet~al., 2021; Wang etalet~al., 2021; Vijaykumar etalet~al., 2023).
Often, at least for simple lens models, it is useful to write the time delay and frequency in dimensionless quantities, which can be defined as (Mishra etalet~al., 2021),
@@
$\begin{split}F(\nu,\boldsymbol{y})&=\frac{\nu}{i}\int d^{2}\boldsymbol{x}\exp[2\pi i\nu\tau_{d}(\boldsymbol{x},\boldsymbol{y})],\\ F(\nu,\boldsymbol{y})|_{\rm geo}&=\sum_{j}\sqrt{|\mu_{j}|}\exp\left(2\pi i\nu\tau_{d}(\boldsymbol{x}_{j},\boldsymbol{y})-i\pi n_{j}\right).\end{split}$ (5)
As we will see below, such a dimensionless form allows us to translate results for one particular lens mass ($M_{z}$) or frequency ($f$) to another.
@@ 3 Calculating $F(f)$
Except for very simple lens models, one cannot obtain an analytic solution for Equation (1) or (5). Hence, for a general lens, we need to rely on numerical integration or some other methods (e.g., Ulmer & Goodman, 1995; Diego etalet~al., 2019; Villarrubia-Rojo etalet~al., 2025). In our current work, we primarily use the method described in Ulmer & Goodman (1995, ‘UG method’ hereafter), which relies on the Fourier transformation and contour integration. First, assuming a constant source position, we take the Fourier transformation of $iF(\nu)/\nu$ in Equation (5) defined as,
$\tilde{F}(\tau^{\prime})\equiv\int d\nu\>\exp(-2\pi i\nu\tau^{\prime})\frac{iF(\nu)}{\nu}.$ (6)
@@
$\tilde{F}(\tau^{\prime})=\int d^{2}\boldsymbol{x}\>\delta[\tau_{d}(\boldsymbol{x};\boldsymbol{y})-\tau^{\prime}].$ (7)
@@
To determine $\tilde{F}(t)$, we use contour integration. As discussed in Mishra etalet~al. (2021), the area between contours of $\tau^{\prime}$ and $\tau^{\prime}+d\tau^{\prime}$ is equal to $\oint dsdl$, where $ds$ is the infinitesimal length along the contour and $dl=d\tau^{\prime}/|\nabla_{\boldsymbol{x}}\tau_{d}(\boldsymbol{x},\boldsymbol{y})|$ is the distance in the orthogonal direction between the two contours. Since there can be more than one contour with a value of $\tau^{\prime}$
@@
Figure 2: Amplification factor, $|F|$, maps for Chang-Refsdal lens with positive values of external shear (i.e., $\gamma>0$). The three rows represent the source plane maps corresponding to $\gamma=0.5/\sqrt{2},0.8/\sqrt{2},0.9/\sqrt{2}$ (i.e., $\mu_{\rm m}=1.33,2.78,5.26$), respectively. In each row, left, middle and right panels correspond to $f=50,100,500$ Hz. In each panel, black solid curves represent the caustic. The black and green stars in the top and middle rows represent the source position for which amplification factor curves are shown in Figure 4.
@@ 4 Point mass lens
For the point mass lens, the amplification factor, Equation (1), has an analytical solution, which is given as,
$\begin{split}F(f,y)=\exp\left\{\frac{\pi\omega}{4}+\frac{i\omega}{2}\left[\ln\left(\frac{\omega}{2}\right)-2\phi_{\rm m}(y)\right]\right\}\\ \Gamma\left(1-\frac{i\omega}{2}\right){}_{1}F_{1}\left(\frac{i\omega}{2},1;\frac{i\omega y^{2}}{2}\right),\end{split}$ (13)
@@
where $\omega\equiv 2\pi\nu$. $\phi_{\rm m}(y)=(x_{\rm m}-y)^{2}/2-\ln(x_{\rm m})$ with $x_{\rm m}=(y+\sqrt{y^{2}+4})/2$ represents the time delay corresponding to the global minima image in geometric optics such that it arrives at $t_{d}=0$. $_{1}F_{1}(a,b,z)$ is the confluent hypergeometric function. In geometric optics limit ($\nu\tau_{d}\gg 1$), the amplification factor can be written as,
$F(f,y)|_{\rm geo}=\sqrt{|\mu_{+}|}-i\sqrt{|\mu_{-}|}\exp[2\pi i\nu\tau_{d}(x,y)],$ (14)
@@
where $\mu_{+}$ and $\mu_{-}$ are the usual magnification factors of minimum and saddle-point images, respectively.
@@
Lensing by a point mass lens can be described by two parameters: redshifted lens mass ($M_{z}$) and source position ($y$). Due to the axial symmetry in the case of a point mass lens, we can expect the interference pattern to show rings of constructive and destructive interference with the distance between consecutive maxima and minima depending on the lens mass and incoming signal frequency (Bulashenko & Ubach, 2022). The amplification factors for a point mass lens with $M_{z}=200~{\rm M_{\odot}}$ are shown in Figure 1. The top row represents the source plane map of the amplification factor ($|F|$) for three different frequencies, $f=50,100,500$ Hz, and the bottom row represents the $F(t)$, $|F|$, and $\theta_{F}$ for three different randomly chosen source positions. These amplification maps can also represent other lens masses or frequencies for the same source position such that they lead to the same value of dimensionless frequency ($\nu=4{\rm G}M_{z}/{\rm c^{3}}$).
@@
In the case of an isolated point mass lens, we always have the formation of two images in geometric optics, with the saddle-point image becoming less and less amplified (approaching zero), and the minima image amplification factor approaches one as the source moves away from the optical axis (i.e., origin). At the same time, the time delay between the two images also increases. Using the geometric optics approximation from Equation (14), we can see that the increase in time delay implies that oscillations in amplification factor would become more rapid at a given frequency, and we move towards geometric optics at relatively lower frequencies. The larger fractional difference in the amplification of two images decreases the amplitude of the oscillations in the amplification factor. Both of the above can be seen from the dashed-dotted curves in the bottom row of Figure 1. In addition, we can also see that the amplification factor under geometric optics approximation and wave optics mainly differs at frequencies lower than the first peak, implying that above the frequency of the first peak, the contribution to the amplification factor is well localized around the stationary points. For the perfect alignment (i.e., $y=0$) case, we have the formation of a ring with zero time delay, implying that we reach geometric optics at $f=\infty$ and wave effects are important for every observable frequency.
@@
Figure 5: Amplification factor, $|F|$, maps for binary lens with $(q_{1},q_{2})=(0.5,0.5)$ and $M_{T}=200~{\rm M_{\odot}}$. In each row, left, middle, and right panels are corresponding to $f=50,100,500$ Hz. From top-to-bottom, we decrease the distance (2LL) between the two-point masses. In each panel, black solid curves represent the caustic structure. The black solid stars are source positions for which we show the amplification factor in Figure 6.
Figure 6: Example amplification factor, $F(f)$, for binary lens case. Source positions are shown by solid stars in the third and fourth rows of Figure 5. The left, middle, and right panels show the F(t)F(t)$|F|$, and $\theta_{F}$ curves, respectively. The solid curves are obtained with the UG method and dashed curves represent the amplification factors obtained under geometric optics. In the bottom panel, for comparison, the dashed-dotted curves represent the amplification factor curves for a point mass lens with $M_{z}=200~{\rm M_{\odot}}$ and $y=|\boldsymbol{y}|$.